OpenStudy (anonymous):

If x-9=2y and x+3=5y what is the value of x

OpenStudy (anonymous):

Well first simplify the equations into Y= form: Y=(X/2) + (3/5) Y=(x/5) + (3/5) Then graph it into a calc ( https://www.desmos.com/calculator) This is the graph: https://www.desmos.com/calculator/4ei5ouebp0 The intersection is the answer! (17,4) This means X=17 and Y=4 Sorry for the late response but I had to graph it!